Radioactive isotopes with high atomic numbers often decay decay to form isotopes that are themselves radioactive, and once form new isotopes. sometimes a series of such decays occurs over many steps until a stable nucleus is formed. The following series of decays occurs: Polonium-218 decays with emission of an alpha particle to form X, which emits a beta particle to form Y, which emits an alpha particle to form Z. Identify X, Y, and Z.
Answer #1

Finding X: Alpha particle has 2 unit positive charge and 4 unit mass number.

Hence when an alpha particle is emitted, the atomic mass decreases by 4 and the atomic number decreases by  2

When Po-218 emits alpha particles, the mass number decreases to 218-4 = 214.

and atomic number decreases from 84(Po) to 84 - 2 = 82 (Pb).

Hence X is Pb-214 or lead-214

Finding Y: Beta particle has unit negative charge and almost no mass

Hence when a beta particle is emitted, the atomic mass remains same and the atomic number increases by 1.

When Pb-214 emits a beta particle, the mass number remains same

and the atomic number increases from 82(Pb) to 83 (Bismuth, Bi)

Hence Y is Bismuth-214 or Bi-214.

Finding Z:

When Bi-214 emits alpha particles, the mass number decreases to 214-4 = 210.

and atomic number decreases from 83(Bi) to 83 - 2 = 81 (Thallium, Tl).

Hence Z is Thallium-210 or Tl-210.